# Assessing Universal Relations for Rapidly Rotating Neutron Stars: Insights from an Interpretable Deep Learning Perspective

### v1

@@ IV.2.3 Relations for the star’s reduced Hexadecapole deformation
Although we employed a high-resolution grid $\mathrm{MDIV}\times\mathrm{SDIV}=751\times 1501$ for stellar models with axial ratios $r_{\mathrm{pole}}/r_{\mathrm{eq}}\in[0.90,0.98]$, the numerical extraction of the mass hexadecapole moment becomes unreliable in most cases for $r_{\mathrm{pole}}/r_{\mathrm{eq}}>0.94$. Given the substantial computational cost associated with the employment of a higher-resolution numerical grid for constructing rapidly rotating NSs in equilibrium, we confine our analysis to stellar models with axial ratios $r_{\mathrm{pole}}/r_{\mathrm{eq}}\leq 0.94$, where the numerical extraction of the hexadecapole remains accurate. This dataset encompasses NS configurations characterized by parameter ranges of $C\in[0.876,3.075]\times 10^{-1}$, $\chi\in[0.247,0.782]$, $\bar{Q}\in[1.313,14.218]$$\bar{S}_{3}\in[1.736,31.034],$ and $\bar{M}_{4}\in[2.655,497.192]$. Above all, it should be highlighted that the relevant dataset consists of $1022628$ stellar models, with $75\%$ allocated for training and the remaining $25\%$ reserved for testing. Table 2 provides the parameter ranges utilized for training and testing for the models associated with the $\bar{M}_{4}$ and $\bar{M}_{4}/\bar{S}_{3}$ parameters.
@@
Figure 23: Relative error distributions on the test dataset for the proposed ANN models described by Eqs. (20) and (21), highlighting their prediction performance.
It is noteworthy that only $1.06\%$ of stellar models exhibit relative deviations larger than $3\%$ for the model (20), while the model (21) shows even better performance, with just $0.6\%$ of stellar models having deviations exceeding $1\%$. For the $\bar{M}_{4}/\bar{S}_{3}$ model, the majority of observed deviations are concentrated in NS configurations with $\bar{Q}\lesssim 7$, while deviations across other parameters appear randomly distributed. In contrast, the model for the $\bar{M}_{4}$ exhibits no discernible pattern in error variance throughout the ($\chi,C,\bar{Q},\bar{S}_{3}$) parameter space. In both model representations, a detailed analysis of error distributions across different EoS categories reveals that the hadronic class exhibits the highest maximum relative deviations. In contrast, the hybrid and hyperonic categories show comparatively lower maximum errors—approximately $5.87\%$ and $5.25\%$ for the $\bar{M}_{4}/\bar{S}_{3}$ regression model (20), and around $5.04\%$ and $4.40\%$ for the $\bar{M}_{4}$ model (21), respectively (see e.g., Appendix C for the corresponding violin plots). Within each EoS class, the EoS models exhibiting the largest deviations are identical and ordered similarly—TM1e, DS(CMF)-2 Hybr, and DS(CMF)-3-for both the $\bar{M}_{4}/\bar{S}_{3}$ and $\bar{M}_{4}$ models.
@@
Figure 24: Predicted $\bar{M}_{4}(C,\bar{Q},\bar{S}_{3})$ data generated by the regression model (21), shown for a discrete set of spin values $\chi\in[0.25,0.70]$. In all cases, the predictions closely match the test data, reinforcing the reliability of the trained ANN model. The vertical color scale encodes the corresponding values of the reduced quadrupole deformation $\bar{Q}$.
@@
Next, we apply the SHAP method to interpret the predictions of the suggested “deep” EoS-insensitive parametrizations related to the $\bar{M}_{4}/\bar{S}_{3}$ and $\bar{M}_{4}$ fits. For the dataset consisting of rapidly rotating NS configurations with axial ratio $r_{\mathrm{pole}}/r_{\mathrm{eq}}\leq 0.94$, we use a subset $X^{{}^{\prime}}_{1}$ which includes $10,000$ NS configurations with parameters $C\in[0.090,0.307],\chi\in[0.249,0.768],\bar{Q}\in[1.317,13.448],\bar{S}_{3}\in[1.742,28.910]$, to provide an interoperability of each model’s predictions. For both models (20) and (21), we assess how variations in the parameter space influence their expected outputs. The corresponding SHAP values, shown in Figures 25 and 26, highlight the contribution of the star’s parameters to the models’ inference.
@@
For the regression model defined in Eq. (20), the SHAP analysis (Fig. 25, top panel) reveals that each global parameter of the NS’s ensemble can exert either a positive or negative influence on the estimator’s mean value. We see that the most (less) compact stars have a negative (positive) contribution. Notably, stars exhibiting greater quadrupole deformation tend to contribute positively to the prediction. Conversely, the most rapidly rotating NSs are associated with negative contributions, while those with lower rotation rates have a positive effect on the inference. The bottom panel of Fig. 25 presents the mean absolute SHAP values for each input feature, quantifying their average impact on the model’s predictions across the $X^{\prime}_{1}$ test subset. The features $\chi$, $C$, and $\bar{Q}$ exhibit mean contributions of $1.20$, $1.11$, and $1.09$, respectively. The comparable magnitude of these contributions suggests a well-balanced sensitivity of the DL model (20) to the input parameters, enabling robust and accurate inference of the $\bar{M}_{4}/\bar{S}_{3}$ across a wide range of stellar configurations.
As shown in the top panel of Fig. 26, the SHAP analysis for the regression model in Eq. (21) demonstrates that each global property of the employed NS ensemble can variably influence the predicted etsimator’s value, with contributions that may be either positive or negative depending on the specific parameter configuration. For both $\chi$ and $\bar{Q}$, feature values at both low and high extremes contribute variably—either positively or negatively—to the model’s predictions. This variability suggests an absence of a clear or consistent SHAP value pattern associated with these parameters in guiding the optimal solution. In contrast, the influence of the $C$ and $\bar{S}_{3}$ parameters is more pronounced. Higher values of both $C$ and $\bar{S}_{3}$ tend to contribute positively to the model’s predictive accuracy, whereas lower values are predominantly associated with negative contributions. The bottom panel of Fig. 26 displays the mean absolute SHAP values for each input feature, providing a quantitative measure of their average influence on the model’s predictions. The input features $\chi$, $C$, $\bar{Q}$, and $\bar{S}_{3}$ exhibit mean contributions of $1.99$, $51.97$, $54.03$, and $136.75$, respectively. While the mean contributions of $C$ and $\bar{Q}$ are comparable, the $\bar{S}_{3}$ parameter stands out with a dominant mean influence, effectively guiding the suggested deep learning model toward robust and accurate inference of $\bar{M}_{4}$ across a the test subsample of stellar configurations.
@@
In this formulation, the regression hypersurface in Eq. (22) accurately reproduces the test data, achieving a maximum relative error of $\leq 7.52\%$, which marks an improvement over the previous model parametrized by Eq. (21) in terms of the maximum fractional deviation metric. Therefore, it demonstrates consistent performance across diverse EoS models, establishing it also as a well-defined EoS-insensitive relation. In contrast to the previous model (with only $0.6\%$ of stellar configurations exceeding $1\%$ deviation), the current simplest one yields $2.38\%$ of NSs with deviations above this threshold. Moreover, as with the previous $\bar{M}_{4}$ DL model, the distribution of errors exhibits no apparent systematic trends, indicating a random variance across the $(C,\bar{Q},\bar{S}_{3})$ parameter space. An assessment of the model’s error distribution across the employed EoS categories reveals that the hadronic class exhibits the largest maximum relative deviation, primarily attributed to the TM1e EoS model. In contrast, the hybrid and hyperonic classes exhibit notably lower relative errors, restricted to approximately $5.56\%$ and $4.32\%$, respectively. These deviations are chiefly driven by the QHC19-D and DS(CMF)-3 EoS models.
It is also insightful to perform a SHAP analysis for the newly proposed $\bar{M}_{4}$ regression model. As illustrated in the top panel of Fig. 27, the SHAP values corresponding to the ANN model (22) reveal that the utilized input parameters of the NS ensemble influence the model’s predictions in varying degrees. These contributions can be either positive or negative, contingent on the NS parameters of interest. For example, we can see that stars with higher compactness tend to exert a negative influence on the prediction, while less compact configurations contribute positively. Notably, stellar configurations exhibiting the largest quadrupole and octupole deformations generally have a positive impact on the estimated value. In contrast, configurations with lower deformation parameters display mixed contributions, which can be either positive or negative depending on the specific parameter combination.
@@
In addition, an examination of the mean absolute SHAP values for the current $\bar{M}_{4}$ trained model shows that the compactness $C$, the reduced spin octupole $\bar{S}_{3}$, and the normalized mass quadrupole $\bar{Q}$ contribute, on average as $37.59$, $25.19$, and $16.19$, respectively, to the model’s output (see e.g., Fig. 27 bottom panel). Unlike the previous $\bar{M}_{4}(\chi,C,\bar{Q},\bar{S}_{3})$ model, compactness $C$ now plays the most significant role. On average, the model’s prediction is strongly influenced by how compact the star is. The second most important feature, $\bar{S}_{3}$, is also closely related to the mass quadrupole moment $\bar{Q}$, reflecting known universal connections between different multipole moments in rotating NSs. In any case, the involvement of all features reflects their collective importance in capturing the underlying universal behavior across different EoSs.
@@ IV.3 Relation for the star’s equatorial radius
$\bar{\mathcal{R}}_{\mathrm{model}}=\hat{\bar{\mathcal{R}}}_{\theta\star}(\bar{M},\chi,\bar{Q},\bar{S}_{3})$ (23)
@@
Table 11: Evaluation measures for the parametrization given by the Eq. (23) regarding both the training and test sets.
@@
Above all, the suggested regression model serves as an invaluable tool for fitting functions that aim to infer the NS surface [151, 150, 153], with direct applications to current observatories like NICER telescope [3, 4, 5] and future missions like the enhanced X-ray Timing and Polarimetry (eXTP) [198], and the Spectroscopic Time-Resolving Observatory for Broadband Energy X-rays (STROBE-X) [199, 200]. These forthcoming missions are expected to surpass the current capabilities of NICER in enhancing the precision of parameter estimation for NS radii.
Finally, for each NS configuration in the test subset $X_{1}$, we analyze the individual contribution of each parameter to the model’s prediction. As illustrated in Fig. 29 (top panel), each feature can exert either a positive or negative influence on the model’s output $\mathbb{E}_{X}(\hat{\bar{\mathcal{R}}}_{\theta^{\star}})$, thus leading the final inferences to accurate predictions. The highest (lowest) values of the parameters $\bar{M},\chi,\bar{Q},\bar{S}_{3}$ have a positive (negative) SHAP contribution.
@@ V Summary and Conclusions
Furthermore, the utilization of the SHapley Additive exPlanations (SHAP) [155, 156, 154] method was of paramount importance in the interpretation and explainability of the predictions generated by the proposed feed-forward network. This framework elucidates the contributions of individual stars’ parameters to the DL model’s output, thereby enhancing the transparency and interpretability of its decisions. More specifically, it enabled a systematic attribution of the influence of each input feature, thereby providing valuable insights into the factors driving the model’s predictions. In addition, this interpretative approach enhances our comprehension of the model’s behavior, corroborates its dependability, and identifies potential avenues for enhancement.
@@
For the spacetime multipole moments, we developed a new EoS-independent relation $\bar{Q}(C,\chi,\mathcal{D})$ (15) with fractional difference better than $2.32\%$, effectively embedding the established $\bar{I}(\bar{Q})$ correlation. More specifically, the suggested model reproduces the expected quasi-universal behavior of the $\bar{Q}(C)$ relation and further supports the robustness of predictions for the newly identified $\bar{Q}(\mathcal{D})$ correlation across the test data, provided that the dimensionless spin parameter $\chi$ is held constant. Above all, SHAP analysis indicates that $\mathcal{D}$ significantly influences the inference, while $C$ and $\chi$ contribute to the model’s ability to capture the underlying universal patterns. For the octupole deformation, we revisited the $\bar{S}_{3}(\log\bar{Q})$ relation (16), achieving fractional difference better than $6.64\%$. In addition, we further introduced an improved new parameterization, $\bar{S}_{3}(\chi,\mathcal{D},\log\bar{Q})$ (17), which demonstrates relative deviation better than $3.32\%$. In addition, the analysis uncovered a robust, previously unidentified universal correlation between $\bar{S}_{3}$ and $\mathcal{D}$ for NS sequences with fixed spin $\chi$ values. The associated trained model (17) successfully captures this EoS-insensitive behavior, underscoring its effectiveness. In any case, within this EoS-insensitive framework, $\mathcal{D}$ and $\log\bar{Q}$ have a pronounced impact on the model’s predictions, while $\chi$ contributes significantly to the identification of spin-related patterns. For the mass hexadecapole, standard relations $\bar{M}_{4}(\log\bar{Q})$ (18) and $(\bar{M}_{4}/\bar{S}_{3})(\log\bar{Q})$ (19) achieved accuracies of $14.34\%$ and $13.22\%$, respectively. The inclusion of higher-order corrections in the $\bar{M}_{4}/\bar{S}_{3}$ and $\bar{M}_{4}$ predictions led to improved accuracies of $9.19\%$ and $8.11\%$, as shown in Eqs. (20) and (21), respectively. In the $(\bar{M}_{4}/\bar{S}_{3})(\chi,C,\bar{Q})$ model, the input features contribute with comparable mean SHAP values, indicating a balanced influence. In contrast, for the $\bar{M}_{4}(\chi,C,\bar{Q},\bar{S}_{3})$ model, $\bar{S}_{3}$ emerges as the most influential feature, followed by moderate contributions from $C$ and $\bar{Q}$, while $\chi$ exhibits the weakest impact. In addition, the marginal contribution of $\chi$ prompted us to explore a simplified $\bar{M}_{4}(C,\bar{Q},\bar{S}_{3})$ model. Despite its reduced parameter space, this formulation achieves a relative error $\leq 7.52\%$, demonstrating satisfactory predictive accuracy. In this model, parameter $C$ emerges as the dominant feature, followed by contributions from $\bar{S}_{3}$ and $\bar{Q}$. While this simpler relation offers a viable and good alternative, the more comprehensive model $\bar{M}_{4}(\chi,C,\bar{Q},\bar{S}_{3})$ remains accurate when spin information is available. Specifically, for the simpler model, $2.38\%$ of NS configurations exhibit relative deviations exceeding $1\%$, whereas this fraction decreases to $0.6\%$ for the more comprehensive formulation.
@@
We also proposed a new EoS-independent relation $\bar{\mathcal{R}}(\bar{M},\chi,\bar{Q},\bar{S}_{3})$ for the equatorial radius (23) with relative deviations better than $3.81\%$. Among the input parameters, $\bar{M}$ and $\bar{Q}$ exert the most significant influence on the model’s inference, while $\chi$ and $\bar{S}_{3}$ play comparatively smaller roles. Mostly, this accurate regression model is a crucial asset for fitting functions focused on inferring NS surfaces [151, 153, 150]. It has direct applications for current observatories such as the NICER telescope [3, 4, 5] and upcoming missions like the enhanced X-ray Timing and Polarimetry (eXTP) [198] and the Spectroscopic Time-Resolving Observatory for Broadband Energy X-rays (STROBE-X) [199, 200]. These upcoming missions are expected to surpass NICER’s current capabilities, providing enhanced precision in measurements of neutron star radii.

### v2

@@ IV.2.3 Relations for the star’s reduced Hexadecapole deformation
Although we employed a high-resolution grid $\mathrm{MDIV}\times\mathrm{SDIV}=751\times 1501$ for stellar models with axial ratios $r_{\mathrm{pole}}/r_{\mathrm{eq}}\in[0.90,0.98]$, the numerical extraction of the mass hexadecapole moment becomes unreliable in most cases for $r_{\mathrm{pole}}/r_{\mathrm{eq}}>0.94$. Given the substantial computational cost associated with the employment of a higher-resolution numerical grid for constructing rapidly rotating NSs in equilibrium, we confine our analysis to stellar models with axial ratios $r_{\mathrm{pole}}/r_{\mathrm{eq}}\leq 0.94$, where the numerical extraction of the hexadecapole remains accurate. Above all, it should be highlighted that the relevant dataset consists of $1022628$ stellar models, with $75\%$ allocated for training and the remaining $25\%$ reserved for testing. Table 8 provides the parameter ranges utilized for training and testing for the models associated with the $\bar{M}_{4}$ and $\bar{M}_{4}/\bar{S}_{3}$ parameters.
Table 8: Parameter ranges of the rotating NS configurations used for training and testing in the $\bar{M}_{4}$ and $\bar{M}_{4}/\bar{S}_{3}$ models.
Feature Training Set Test Set
min max min max
$C\times 10^{-1}$ 0.876 3.075 0.882 3.075
$\chi\times 10^{-1}$ 2.471 7.818 2.472 7.803
$\bar{Q}$ 1.313 14.218 1.314 14.115
$\bar{S_{3}}$ 1.736 31.034 1.738 30.799
$\bar{M_{4}}$ 2.655 497.192 2.658 491.585
$\bar{M_{4}}/\bar{S}_{3}$ 1.525 16.021 1.524 15.961
@@
Figure 24: Relative error distributions on the test dataset for the proposed ANN models described by Eqs. (22) and (23), highlighting their performance.
It is noteworthy that only $1.06\%$ of stellar models exhibit relative deviations larger than $3\%$ for the model (22), while the model (23) shows even better performance, with just $0.6\%$ of stellar models having deviations exceeding $1\%$. For the $\bar{M}_{4}/\bar{S}_{3}$ model, the majority of observed deviations are primarily found in NS configurations with $\bar{Q}\lesssim 7$, corresponding to the most compact stellar models with $C\in[0.136,0.256]$, independent of the spin. In contrast, the model for the $\bar{M}_{4}$ exhibits no discernible pattern in error variance throughout the ($\chi,C,\bar{Q},\bar{S}_{3}$) parameter space. In both model representations, a detailed analysis of error distributions across different EoS categories reveals that the hadronic class exhibits the highest maximum relative deviations. In contrast, the hybrid and hyperonic categories show comparatively lower maximum errors—approximately $5.87\%$ and $5.25\%$ for the $\bar{M}_{4}/\bar{S}_{3}$ regression model (22), and around $5.04\%$ and $4.40\%$ for the $\bar{M}_{4}$ model (23), respectively (see e.g., Appendix C for the corresponding violin plots). Within each EoS class, the EoS models exhibiting the largest deviations are identical and ordered similarly—TM1e, DS(CMF)-2 Hybr, and DS(CMF)-3-for both the $\bar{M}_{4}/\bar{S}_{3}$ and $\bar{M}_{4}$ models.
@@
Figure 25: $\bar{M}_{4}-C$ plane: Predicted $\bar{M}_{4}(C,\bar{Q},\bar{S}_{3})$ data generated by the regression model (23), shown for a discrete set of spin values $\chi\in[0.25,0.70]$. Data points are color-coded by the ANN model’s absolute relative error [%], while the red star marks the configuration with the maximum relative error. In all cases, the predictions closely match the illustrated test data, reinforcing the reliability of the trained ANN model.
@@
Next, we apply the SHAP method to interpret the predictions of the suggested “deep” EoS-insensitive parametrizations related to the $\bar{M}_{4}/\bar{S}_{3}$ and $\bar{M}_{4}$ fits. For the dataset consisting of rapidly rotating NS configurations with axial ratio $r_{\mathrm{pole}}/r_{\mathrm{eq}}\leq 0.94$, we use a subset $X^{{}^{\prime}}_{1}$ which includes $10,000$ NS configurations with parameters $C\in[0.090,0.307],\chi\in[0.249,0.768],\bar{Q}\in[1.317,13.448],\bar{S}_{3}\in[1.742,28.910]$, to provide interpretability of each model’s predictions. For both models (22) and (23), we assess how the chosen parameters can impact the associated expected outputs. The corresponding SHAP values, shown in Figures 26 and 27, highlight the contribution of the star’s parameters to the models’ inference.
@@
For the regression model defined in Eq. (22), the SHAP analysis (Fig. 26, top panel) reveals that each global parameter of the NS’s ensemble can exert either a positive or negative influence on the estimator’s mean value. We see that the most (less) compact stars have a negative (positive) contribution. Notably, stars exhibiting larger quadrupole deformation tend to contribute positively relative to the model’s expected output. Conversely, the most rapidly rotating NSs are associated with negative SHAP attributions, whereas lower-spin configurations show positive ones. The bottom panel of Fig. 26 presents the mean absolute SHAP values for each input feature, quantifying their average impact across the $X^{\prime}_{1}$ test subset. The features $\chi$, $C$, and $\bar{Q}$ exhibit mean contributions of $1.20$, $1.11$, and $1.09$, respectively. In this case, the average SHAP contributions of the physical features are found to be comparable. As shown in previous parametrizations, this is not always the case, but it can naturally occur when correlations within the universal parametrization, together with the way the trained model’s weights distribute importance across the inputs, lead to a more balanced SHAP contribution among features. Nevertheless, the model’s input parameters enable robust and accurate inference of $\bar{M}_{4}/\bar{S}_{3}$ across a wide range of stellar configurations.
As shown in the top panel of Fig. 27, the SHAP analysis for the regression model in Eq. (23) demonstrates that each global parameter of the employed NS ensemble can variably influence the predicted expected value $\mathbb{E}_{X}(\hat{\mathcal{M}}^{\prime}_{\theta^{\star}})$, with contributions that may be either positive or negative depending on the specific parameter configuration. For both $\chi$ and $\bar{Q}$, low and high extreme values exhibit mixed SHAP attributions—positive and negative—relative to the model’s expected value prediction. This variability suggests an absence of a clear or consistent SHAP value pattern associated with these parameters. In contrast, the influence of the $C$ and $\bar{S}_{3}$ parameters is more pronounced. Higher values of both $C$ and $\bar{S}_{3}$ tend to contribute positively, whereas lower values are predominantly associated with negative contributions. The bottom panel of Fig. 27 displays the mean absolute SHAP values for each input feature, providing a quantitative measure of their average influence on the model’s predictions. The input features $\chi$, $C$, $\bar{Q}$, and $\bar{S}_{3}$ exhibit mean contributions of $1.99$, $51.97$, $54.03$, and $136.75$, respectively. Above all, while the mean contributions of $C$ and $\bar{Q}$ are comparable, the $\bar{S}_{3}$ parameter stands out with a dominant influence on the model’s expected value $\mathbb{E}_{X}(\hat{\mathcal{M}}^{\prime}_{\theta^{\star}})$.
@@
In this formulation, the regression hypersurface in Eq. (24) accurately reproduces the test data, achieving a maximum relative error of $\leq 7.52\%$. This represents an improvement with respect to the maximum fractional deviation evaluation measure compared to the model in Eq. (23), while maintaining a simpler parametrization. Therefore, it demonstrates consistent performance across diverse EoS models, establishing it also as a well-defined EoS-insensitive relation. In contrast to the previous model (with only $0.6\%$ of stellar configurations exceeding $1\%$ deviation), the current simplest one yields $2.38\%$ of NSs with deviations above this threshold. Moreover, as with the previous $\bar{M}_{4}$ DL model, the distribution of errors exhibits no apparent systematic trends, indicating a random variance across the $(C,\bar{Q},\bar{S}_{3})$ parameter space. An assessment of the model’s error distribution across the employed EoS categories reveals that the hadronic class exhibits the largest maximum relative deviation, primarily attributed to the TM1e EoS model. In contrast, the hybrid and hyperonic classes exhibit notably lower relative errors, restricted to approximately $5.56\%$ and $4.32\%$, respectively. These deviations are chiefly driven by the QHC19-D and DS(CMF)-3 EoS models.
It is also insightful to perform a SHAP analysis for the newly proposed $\bar{M}_{4}$ regression model. As illustrated in the top panel of Fig. 28, the SHAP values corresponding to the ANN model (24) reveal that the utilized input parameters of the NS ensemble influence the model’s predictions in varying degrees. These contributions can be either positive or negative, contingent on the NS parameters of interest. For example, we can see that stars with higher compactness tend to exert a negative influence on the model’s expected value, while less compact configurations contribute positively. Notably, stellar configurations exhibiting the largest quadrupole and octupole deformations generally have a positive impact. In contrast, configurations with lower deformation parameters display mixed contributions, which can be either positive or negative depending on the specific parameter combination.
@@
In addition, an examination of the mean absolute SHAP values for the current $\bar{M}_{4}$ trained model shows that the compactness $C$, the reduced spin octupole $\bar{S}_{3}$, and the normalized mass quadrupole $\bar{Q}$ contribute, on average as $37.59$, $25.19$, and $16.19$, respectively (see e.g., Fig. 28 bottom panel). Unlike the previous $\bar{M}_{4}(\chi,C,\bar{Q},\bar{S}_{3})$ model, compactness $C$ now plays the most significant role. On average, the model’s prediction is strongly influenced by how compact the star is. The second most important feature, $\bar{S}_{3}$, is also closely related to the mass quadrupole moment $\bar{Q}$, reflecting known universal connections between different multipole moments in rotating NSs. In any case, the involvement of all features reflects their collective importance in capturing the underlying universal behavior across different EoSs.
@@ IV.3 Relation for the star’s equatorial radius
$\bar{\mathcal{R}}_{\mathrm{model}}=\hat{\bar{\mathcal{R}}}_{\theta^{\star}}(\bar{M},\chi,\bar{Q},\bar{S}_{3})$ (25)
@@
Above all, the suggested regression model serves as an invaluable tool for fitting functions that aim to infer the NS surface [151, 150, 153], where $R_{\mathrm{eq}}$ serves as an input parameter, with direct applications to current observatories like NICER telescope [3, 4, 5] and future missions like the enhanced X-ray Timing and Polarimetry (eXTP) [199], and the Spectroscopic Time-Resolving Observatory for Broadband Energy X-rays (STROBE-X) [200, 201]. These forthcoming missions are expected to surpass the current capabilities of NICER in enhancing the precision of parameter estimation for NS radii.
Finally, for each NS configuration in the test subset $X_{1}$, we analyze the individual contribution of each parameter to the model’s prediction. As illustrated in Fig. 30 (top panel), each feature can exert either a positive or negative influence on the model’s expected value $\mathbb{E}_{X}(\hat{\bar{\mathcal{R}}}_{\theta^{\star}})$. The highest (lowest) values of the parameters $\bar{M},\chi,\bar{Q},\bar{S}_{3}$ have a positive (negative) SHAP contribution.
@@ V Summary and Conclusions
Furthermore, the utilization of the SHapley Additive exPlanations (SHAP) [155, 156, 154] method was of paramount importance in the interpretation and explainability of the predictions generated by the proposed feed-forward network. This framework elucidates the contributions of individual stars’ parameters to the DL model’s expected value, thereby enhancing the transparency and interpretability of its decisions. More specifically, it enabled a systematic attribution of the influence of each input feature, thereby providing valuable insights into the factors driving the model’s predictions. In addition, this interpretative approach enhances our comprehension of the model’s behavior, corroborates its dependability, and identifies potential avenues for enhancement.
@@
For the spacetime multipole moments, we developed a new EoS-independent relation $\bar{Q}(C,\chi,\mathcal{D})$ (17) with a fractional difference better than $2.32\%$, effectively embedding the established $\bar{I}(\bar{Q})$ correlation. More specifically, the suggested model reproduces the expected quasi-universal behavior of the $\bar{Q}(C)$ relation and further supports the robustness of predictions for the newly identified $\bar{Q}(\mathcal{D})$ correlation across the test data, provided that the dimensionless spin parameter $\chi$ is held constant. Above all, SHAP analysis indicates that $\mathcal{D}$ significantly influences the model’s expected value, while $C$ and $\chi$ contribute to the model’s ability to capture the underlying universal patterns. For the octupole deformation, we revisited the $\bar{S}_{3}(\log\bar{Q})$ relation (18), achieving fractional difference better than $6.64\%$. In addition, we further introduced an improved new parameterization, $\bar{S}_{3}(\chi,\mathcal{D},\log\bar{Q})$ (19), which demonstrates relative deviation better than $3.32\%$. In addition, the analysis uncovered a robust, previously unidentified universal correlation between $\bar{S}_{3}$ and $\mathcal{D}$ for NS sequences with fixed spin $\chi$ values. The associated trained model (19) successfully captures this EoS-insensitive behavior, underscoring its effectiveness. In any case, within this EoS-insensitive framework, $\mathcal{D}$ and $\log\bar{Q}$ have a pronounced impact on the model’s predictions, while $\chi$ contributes significantly to the identification of spin-related patterns. For the mass hexadecapole, standard relations $\bar{M}_{4}(\log\bar{Q})$ (20) and $(\bar{M}_{4}/\bar{S}_{3})(\log\bar{Q})$ (21) achieved accuracies of $14.34\%$ and $13.22\%$, respectively. The inclusion of higher-order corrections in the $\bar{M}_{4}/\bar{S}_{3}$ and $\bar{M}_{4}$ predictions led to improved accuracies of $9.19\%$ and $8.11\%$, as shown in Eqs. (22) and (23), respectively. In the $(\bar{M}_{4}/\bar{S}_{3})(\chi,C,\bar{Q})$ model, the input features contribute with comparable mean SHAP values, indicating a balanced influence. In contrast, for the $\bar{M}_{4}(\chi,C,\bar{Q},\bar{S}_{3})$ model, $\bar{S}_{3}$ emerges as the most influential feature, followed by moderate contributions from $C$ and $\bar{Q}$, while $\chi$ exhibits the weakest impact. In addition, the marginal contribution of $\chi$ prompted us to explore a simplified $\bar{M}_{4}(C,\bar{Q},\bar{S}_{3})$ model. Despite its reduced parameter space, this formulation achieves a relative error $\leq 7.52\%$, demonstrating satisfactory predictive accuracy. In this model, parameter $C$ emerges as the dominant feature, followed by contributions from $\bar{S}_{3}$ and $\bar{Q}$. While this simpler relation offers a viable and good alternative, the more comprehensive model $\bar{M}_{4}(\chi,C,\bar{Q},\bar{S}_{3})$ remains accurate when spin information is available. Specifically, for the simpler model, $2.38\%$ of NS configurations exhibit relative deviations exceeding $1\%$, whereas this fraction decreases to $0.6\%$ for the more comprehensive formulation.
@@
We also proposed a new EoS-independent relation $\bar{\mathcal{R}}(\bar{M},\chi,\bar{Q},\bar{S}_{3})$ for the equatorial radius (25) with relative deviations better than $3.81\%$. Among the input parameters, $\bar{M}$ and $\bar{Q}$ exert the most significant contribution on the model’s expected value, while $\chi$ and $\bar{S}_{3}$ play comparatively smaller roles. Mostly, this accurate regression model is a crucial asset for fitting functions focused on inferring NS surfaces [151, 153, 150]. It has direct applications for current observatories such as the NICER telescope [3, 4, 5] and upcoming missions like the enhanced X-ray Timing and Polarimetry (eXTP) [199] and the Spectroscopic Time-Resolving Observatory for Broadband Energy X-rays (STROBE-X) [200, 201]. These upcoming missions are expected to surpass NICER’s current capabilities, providing enhanced precision in measurements of neutron star radii.